The density of mercury is 13.6 g/cm3 . What is the mass of 6.50 cm3 of mercury?


1
Expert's answer
2021-02-01T04:03:42-0500

m = V*ρ = 6.5*13.6 = 88.4 g

Answer: 88.4 g